Hello, My name is Kescha and I was trying to work out a problem and it's solve the equation 20-3x= 15+2x
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

20-3x= 15+2x
Add 3x to both sides
20 = 15 + 5x
Subtract 15
5 = 5x
Divide by 5
x = 1

Answer by stanbon(75887)

20-3x= 15+2x
---
Add 3x to both sides.
Subtract 15 from both sides.
----
5 = 5x
Divide both sides by 5 to get:
x = 1
=====================
Cheers,
Stan H.
=====================